OpenStudy (anonymous):

if this is true prove it, if false make a counter z bar = mod(z) ^ 2 / z, where z can't equal 0

OpenStudy (anonymous):

\[\overline{z}=\frac{|z|^2}{z}\] If \(z=x+iy\), then \[\overline{z}=x-iy\\ |z|=\sqrt{x^2+y^2}\] So you have \[\begin{align*}x-iy&=\frac{x^2+y^2}{x+iy}\\ (x-iy)&=\frac{x^2+y^2}{x+iy}\cdot\frac{x-iy}{x-iy}\\ (x-iy)&=\frac{x^2+y^2}{x^2-i^2y^2}(x-iy)\\ (x-iy)&=\color{red}{\frac{x^2+y^2}{x^2+y^2}}(x-iy) \end{align*}\] Clearly, the red part is 1, so the statement is true.

OpenStudy (anonymous):

thank you this was so helpful